OpenStudy (anonymous):

2. Mr. Cooper���������������������������s company rents storage containers to customers. The customers can choose which size fits their needs and Cooper Storage delivers the containers. After the customers fill the containers, the company transports them to a warehouse to be stored until the customer needs the contents. (a) The volume for one size of the storage containers is 574 ft3. The container has a length of 10 ft and a width of 7 ft. How tall would a stack of 6 containers be if they were stacked one on top of another? (b) There are two other different sizes of containers.

OpenStudy (anonymous):

(b) There are two other different sizes of containers. The price for renting a container per month for a residence within 30 miles is: 7’ container: 7 ft by 7 ft by 8 ft $99.00 12’ container: 12 ft by 8 ft by 8 ft $199.00 What is the price per cubic foot for each container? Round prices to the nearest cent.

OpenStudy (blurbendy):

is this multiple choice?

OpenStudy (anonymous):

no, it is a "fill in the blank" with no choises

OpenStudy (anonymous):

and no blank

OpenStudy (anonymous):

i just have to answer it, no choises to choose from

OpenStudy (blurbendy):

for b) 7 x 7 x 8 = 392 392 / 99 = $3.96 can you figure out the 12' container?

OpenStudy (anonymous):

um....no

OpenStudy (blurbendy):

do the same thing I did for the 7' container

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12 * 12 * 12 = 1728

OpenStudy (anonymous):

opps wrong thing

OpenStudy (anonymous):

12 * 8 * 8 = 768

OpenStudy (blurbendy):

yup, then divide it by 199

OpenStudy (anonymous):

768 / 199 = 3.85929648241206

OpenStudy (blurbendy):

yes, or 3.86

OpenStudy (blurbendy):

if you round it

OpenStudy (anonymous):

ok

OpenStudy (blurbendy):

for a) did you mean 10 1/4 feet?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

k 574 = 10.25 * 7 *h solve for h

OpenStudy (anonymous):

um.......392

OpenStudy (blurbendy):

no first multiply 10.25 * 7 = 71.75 now divide both sides by 71.75 h = 574/71.75 = ???

OpenStudy (anonymous):

0.125

OpenStudy (blurbendy):

no 574/71.75 = 8

OpenStudy (blurbendy):

each container is 8 feet and you stack 6 of them so 8 * 6 = ???

OpenStudy (anonymous):

opps....i did it backward

OpenStudy (anonymous):

48

OpenStudy (blurbendy):

good

OpenStudy (anonymous):

thank you now i have to open up a new question....unless youll let me ask one here

OpenStudy (blurbendy):

you can ask it here

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Riley finds out that Cooper’s Storage and Shipping Company is working with a local business to package some office supplies. Some of the supplies are packed inside a cube-shaped box with side lengths of 4 in. These boxes are then packed into a shipping box with dimensions of 18 in. × 9 in. × 4 in. How many boxes of office supplies can be packed into the larger box for shipping? Show your work.

OpenStudy (blurbendy):

18 x 9 x 4 = ???

OpenStudy (blurbendy):

hello?

OpenStudy (anonymous):

soory

OpenStudy (anonymous):

648

OpenStudy (anonymous):

still there?

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

ok so

OpenStudy (blurbendy):

the small cube is 4 * 4 * 4 = 64

OpenStudy (blurbendy):

so what's 648 / 64

OpenStudy (anonymous):

10.125

OpenStudy (blurbendy):

are you sure you wrote down the numbers correctly?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

did you mean the side lengths are 4 1/2 inches?

OpenStudy (anonymous):

i think so...but thats not what i got....even using the calculator to check myself....

OpenStudy (blurbendy):

well if the question is asking 4 1/2 inches then 4.5 * 4.5 * 4.5 = 91.125 18 * 9 * 4.5 = 729 729 / 91.125 = 8

OpenStudy (anonymous):

ok i have that and is that all if it is then im copy-pasting the next question

OpenStudy (blurbendy):

that's all the question seems to be asking, so yes

OpenStudy (anonymous):

3. One of the offices at the company was being repainted. Riley’s father challenged him to figure out how many quarts of paint would be needed. The dimensions of the office are 16 ft by 12 ft by 10 ft. The office floor, ceiling, and door would not be painted and there are no windows in this office. The door opening measures 3 ft by 8 ft. One quart of paint covers 100 ft2. How many quarts of paint would be needed to paint the room, if only full quarts can be purchased? Show your work.

OpenStudy (blurbendy):

door area = 8 * 3 = 24 wall area = 16 * 12 * 10 = ???

OpenStudy (anonymous):

1920

OpenStudy (blurbendy):

wait I did that wrong, one sec

OpenStudy (blurbendy):

Im not sure how many walls there are in the office...

OpenStudy (blurbendy):

4 im guessing

OpenStudy (blurbendy):

so, if there's 4 walls then you need to solve ok so you have 4 walls so you need to solve 2*(12*10) + 2*(16*10)

OpenStudy (anonymous):

120........160......both go to....240........320

OpenStudy (anonymous):

forgot to add....

OpenStudy (anonymous):

560

OpenStudy (blurbendy):

very good. now subtract the door area because it isn't going to be painted

OpenStudy (anonymous):

536...right?

OpenStudy (blurbendy):

yup, now each paint can gives you 100 ft^2 so divide 536/100

OpenStudy (anonymous):

5.36

OpenStudy (blurbendy):

yes, but you can only buy full cans, so you would need 6 cans total to paint the walls

OpenStudy (blurbendy):

are you in an honors class or something? cause these questions are hard lol

OpenStudy (anonymous):

ok let me put the answers in the doccument and then i am done for the day......wait, i still have a assessment

OpenStudy (blurbendy):

oh

OpenStudy (anonymous):

they are hard cas i have a 7th grade math levl

OpenStudy (blurbendy):

ohhhhh, no wonder. dang you're smart

OpenStudy (anonymous):

and im only in 6th grade

OpenStudy (anonymous):

with a 11th grade reading levl